A helium-filled balloon at 305.0 K and 1 atm, contains 0.35 g He, and has a volume of 1.31 L. It is placed in a freezer (T = 240.0 K), and its volume decreases to 0.90 L. Find ΔE for the gas (in joules). (Cp of He = 20.8 J/mol K.)
Answer #1

q = nCΔT
(moles x heat capacity x (Final Temperature - Initial Temperature)

w = -PΔ V
(-pressure x (Final Volume - Initial Volume)

ΔE = q + w

So...

moles of he = 0.35 g /4 g/mol = 0.0875 moles

q = 0.0875 mol x 20.8 J/°mol k x (-65.0 °k) = 45601.92 = -118.3 J

w = -1atm x (0.9 L - 1.31 L) (101.3 J/L atm)<----- conversion factor in atm pressure
W = 41.533 J

ΔE = 41.533 J + (-118.3 J) = -76.767 J ( - sing show energy out from system )